OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)?

OpenStudy (anonymous):

I really need help for this question

OpenStudy (anonymous):

these are the mutiple choice answers btw x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

OpenStudy (anonymous):

y=(-1/3)+3

OpenStudy (anonymous):

oh one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no prob